# TolerantECG: A Foundation Model for Imperfect Electrocardiogram

### v3

@@ 3. Method
3.1. Cardio Feature Retrieval (CFR)
Figure 1. The Cardio Feature Retrieval (CFR) pipeline employs information gathered from Life in the Fastlane website to retrieve relevant waveform criteria for each cardiac disease. Diagnoses are encoded into vector embeddings and stored in a vector database.
Yu et al. (2024) and Yu et al. (2023) employs a ChatGPT-based Retrieval-Augmented Generatio (RAG) system to extract waveform criteria of various cardiac diseases from provided textbooks as external knowledge. However, relying on OpenAI’s API for text embeddings and tokens can become costly for large datasets, and ChatGPT’s reliability, particularly in the medical domain, is not guaranteed (Shen et al., 2023; Cappellani et al., 2024). To address these challenges, we develop Cardio Feature Retrieval (CFR), a knowledge retrieval system that can directly retrieve text description of waveform features associated with a given diagnosis without the need for a Large Language Model (LLM). The retrieved waveform features, along with given metadata from ECG recordings including patient’s gender and age, are used to construct a detailed report description for each ECG signal. The constructed text description is used during contrastive learning to enhance the semantic representations of ECG signals. The schematic of the CFR method is depicted in Figure 1.
@@ 3.2.1. Report Alignment (ReportAlign)
In addition, to extract meaningful features from the report description, we employ BioLinkBERT (Yasunaga et al., 2022) as the Text encoder. BioLinkBERT is a BERT-based model (Devlin et al., 2019) pretrained on extensive biomedical and clinical text datasets from PubMed. The CLS token from the final layer serves as the embedding for the entire text input sequence. We set random initial weights for the ECG encoder and initialize the Text encoder with the pretrained weights. During pretraining, the Text encoder is kept frozen (Liu et al., 2024; Yu et al., 2024; Zhao et al., 2024).
@@
$\mathcal{L}_{\text{ReportAlign}}=-\frac{1}{2}\Bigg{(}$ $\underbrace{\sum_{i=1}^{B}\log\frac{\exp(S_{i}^{\top}T_{i}/\tau)}{\sum_{j=1}^{B}\exp(S_{i}^{\top}T_{j}/\tau)}}_{\text{ecg-to-text}}$
$+$ $\underbrace{\sum_{i=1}^{B}\log\frac{\exp(T_{i}^{\top}S_{i}/\tau)}{\sum_{j=1}^{B}\exp(T_{i}^{\top}S_{j}/\tau)}}_{\text{text-to-ecg}}\Bigg{)},$
where $S_{i}$ and $T_{i}$ are the projected embeddings from the ECG encoder and Text encoder for the $i$-th signal-text pair, $B$ is the training batch size, and $\tau$ is the logit temperature scaling factor.
@@ 3.2.2. Self-supervised learning with dual-mode distillation (DuoDistill)
$\mathcal{L}_{\text{DuoDistill}}=\mathbb{E}\Bigg{[}\sum_{i=1}^{K}\sum_{j=1\text{, }j\neq i}^{K+N}\sum_{d=1}^{D}-\text{softmax}\left(\frac{\mathbf{q}^{i}_{d}-\mathbf{c}_{d}}{\tau_{t}}\right)$
$\cdot\log\Bigg{(}\text{softmax}\left(\frac{\mathbf{p}^{j}_{d}}{\tau_{s}}\right)\Bigg{)}\Bigg{]},$
@@
$\mathcal{L}_{\text{Total}}=\alpha\mathcal{L}_{\text{ReportAlign}}+\beta\mathcal{L}_{\text{DuoDistill}},$
@@ 3.2.3. Alternating training
At the odd steps, the ECG signals are corrupted with up to all three types of noise. The process follows the same flow as the lead-missing branch; however, in this case, the noise expert teacher is updated after backpropagation. This alternating strategy resembles a student progressively learning from one subject to another.
@@ 5. Conclusion
This paper introduces TolerantECG, a novel training approach for an ECG foundation model capable of learning robust ECG representations from noisy and/or lead-missing ECG recordings. The method combines contrastive learning guided by detailed cardio text reports and self-supervised learning with dual-mode distillation that handles the two types of imperfect ECG conditions. Extensive experiments provide a comprehensive analysis of the model’s downstream transferability across all class levels (i.e. downstream tasks) of the PTB-XL dataset. Results demonstrate that the proposed method consistently ranks as the best or second-best performer across all the downstream tasks under various ECG signal conditions. An ablation study highlights that the dual-mode distillation module contributes significantly more to overall performance than the report-alignment contrastive learning module in the TolerantECG framework.

### v4

@@ 3. Method
3.1. Cardiac Feature Retrieval (CFR)
Figure 1. The Cardiac Feature Retrieval (CFR) pipeline employs information gathered from Life in the Fastlane website to retrieve relevant waveform criteria for each cardiac disease. Diagnoses are encoded into vector embeddings and stored in a vector database.
Yu et al. (2024) and Yu et al. (2023) employs a ChatGPT-based Retrieval-Augmented Generatio (RAG) system to extract waveform criteria of various cardiac diseases from provided textbooks as external knowledge. However, relying on OpenAI’s API for text embeddings and tokens can become costly for large datasets, and ChatGPT’s reliability, particularly in the medical domain, is not guaranteed (Shen et al., 2023; Cappellani et al., 2024). To address these challenges, we develop Cardiac Feature Retrieval (CFR), a knowledge retrieval system that can directly retrieve text description of waveform features associated with a given diagnosis without the need for a Large Language Model (LLM). The retrieved waveform features, along with given metadata from ECG recordings including patient’s gender and age, are used to construct a detailed report description for each ECG signal. The constructed text description is used during contrastive learning to enhance the semantic representations of ECG signals. The schematic of the CFR method is depicted in Figure 1.
@@ 3.2.1. Report Alignment (ReportAlign)
In addition, to extract meaningful features from the report description, we employ BioLinkBERT (Yasunaga et al., 2022) as the Text encoder. BioLinkBERT is a BERT-based model (Devlin et al., 2019) pretrained on extensive biomedical and clinical text datasets from PubMed. The CLS token from the final layer serves as the embedding for the entire text input sequence. We set random initial weights for the ECG encoder and initialize the Text encoder with the pretrained weights. During pretraining, the Text encoder is kept frozen (Li et al., 2023; Liu et al., 2024; Yu et al., 2024; Zhao et al., 2024).
@@
$\mathcal{L}_{\text{ReportAlign}}(\mathbf{S},\mathbf{T})=-\frac{1}{2}\Bigg(\underbrace{\sum_{i=1}^{B}\log\frac{\exp(\mathbf{S}_{i}^{\top}\mathbf{T}_{i}/\tau)}{\sum_{j=1}^{B}\exp(\mathbf{S}_{i}^{\top}\mathbf{T}_{j}/\tau)}}_{\text{ecg-to-text}}+\underbrace{\sum_{i=1}^{B}\log\frac{\exp(\mathbf{T}_{i}^{\top}\mathbf{S}_{i}/\tau)}{\sum_{j=1}^{B}\exp(\mathbf{T}_{i}^{\top}\mathbf{S}_{j}/\tau)}}_{\text{text-to-ecg}}\Bigg)$
where $\mathbf{S}_{i}$ and $\mathbf{T}_{i}$ are the projected embeddings from the ECG encoder and Text encoder for the $i$-th signal-text pair, $B$ is the training batch size, and $\tau$ is the logit temperature scaling factor.
@@ 3.2.2. Self-supervised learning with dual-mode distillation (DuoDistill)
$\mathcal{L}_{\text{DuoDistill}}(\mathbf{q},\mathbf{p})=-\mathbb{E}\Bigg[\sum_{i=1}^{K}\sum_{j=1\text{, }j\neq i}^{K+N}\sum_{d=1}^{D}\text{softmax}\left(\frac{\mathbf{q}^{i}_{d}-\mathbf{c}_{d}}{\tau_{t}}\right)\cdot\log\Bigg(\text{softmax}\left(\frac{\mathbf{p}^{j}_{d}}{\tau_{s}}\right)\Bigg)\Bigg]$
@@
$\mathcal{L}_{\text{Total}}=\alpha\mathcal{L}_{\text{ReportAlign}}(\mathbf{S},\mathbf{T})+\beta\mathcal{L}_{\text{DuoDistill}}(\mathbf{q},\mathbf{p}),$
@@ 3.2.3. Alternating training
At the odd steps, the ECG signals are corrupted with up to all three types of noise. The process follows the same flow as the lead-missing branch; however, in this case, the noise expert teacher is updated after backpropagation. This alternating strategy resembles a student progressively learning from one subject to another. The entire training procedure is outlined in Algorithm 1.
Input: ECG signal $E$, report description $R$; mask teacher $\overline{\textit{ECG\_Encoder}}$, noisy teacher $\overline{\overline{\textit{ECG\_Encoder}}}$, student model ECG_Encoder, text model Text_Encoder, total steps $S$
1
2for $s=0$ to S1S-1 do
3   $\mathbf{S}=\textit{ECG\_Encoder}(E)$ // ECG embedding
4   $\mathbf{T}=\textit{Text\_Encoder}(R)$ // Text embedding
5
6  if $s$ is even then
7     $(E_{\text{major}},E_{\text{minor}})=\text{add\_mask}(E)$
8     $\mathbf{q}=\overline{\textit{ECG\_Encoder}}(E_{\text{minor}})$
9     $\mathbf{p}=\textit{ECG\_Encoder}([E_{\text{minor}},E_{\text{major}}])$
10     else
11       $(E_{\text{major}},E_{\text{minor}})=\text{add\_noise}(E)$
12       $\mathbf{q}=\overline{\overline{\textit{ECG\_Encoder}}}(E_{\text{minor}})$
13       $\mathbf{p}=\textit{ECG\_Encoder}([E_{\text{minor}},E_{\text{major}}])$
14       end if
15
16      $\mathcal{L}_{\text{Total}}=\alpha\mathcal{L}_{\text{ReportAlign}}(\mathbf{S},\mathbf{T})+\beta\mathcal{L}_{\text{DuoDistill}}(\mathbf{q},\mathbf{p})$
17
18      Backpropagate $\mathcal{L_{\text{Total}}}$
19
20      if $s$ is even then
21         update($\overline{\textit{ECG\_Encoder}}\leftarrow\textit{ECG\_Encoder}$)
22         else
23           update($\overline{\overline{\textit{ECG\_Encoder}}}\leftarrow\textit{ECG\_Encoder}$)
24           end if
25
26           end for
Algorithm 1 Training process of TolerantECG
@@ 5. Conclusion
This paper introduces TolerantECG, a novel training approach for an ECG foundation model capable of learning robust ECG representations from noisy and/or lead-missing ECG recordings. The method combines contrastive learning guided by detailed cardiac text reports and self-supervised learning with dual-mode distillation that handles the two types of imperfect ECG conditions. Extensive experiments provide a comprehensive analysis of the model’s downstream transferability across all class levels (i.e. downstream tasks) of the PTB-XL dataset. Results demonstrate that the proposed method consistently ranks as the best or second-best performer across all the downstream tasks under various ECG signal conditions. An ablation study highlights that the dual-mode distillation module contributes significantly more to overall performance than the report-alignment contrastive learning module in the TolerantECG framework.